three times a number added to twice a smaller number is 4. twice the smaller number less than twice the larger number is 6. find the numbers


help :(

Answer by rothauserc(4718)

we are given
3i + 2j = 4
and 2j - 2i = 6
therefore 2j = 2i + 6 and
j = i + 3
substitute in the first equation
3i + 2 * (i+3) = 4
3i + 2i + 6 = 4
5i = -2
i = -2/5 and -6/5 + 2j = 4, -6 + 10j = 20, j = 26/10 = 13/5
i = -2/5 and j=13/5